OpenStudy (anonymous):

Find the inverse matrix of: {-4x+y=5 {x-3y=4

OpenStudy (accessdenied):

How would you set up the matrix itself?

OpenStudy (anonymous):

[-4 1 | 5] [1 -3 | 4]

OpenStudy (accessdenied):

not quite sure you can find the inverse of a matrix like that... i think it would be something like this [-4 1][x] [5] [1 -3][y] = [4]

OpenStudy (anonymous):

How do you find x and y?

OpenStudy (accessdenied):

it isn't really asking for x and y, it only wanted the inverse matrix... but now I'm not quite sure how you could figure it out with my way either. :x Sorry, might have to find somebody else, this isn't my strongpoint. :(

OpenStudy (bahrom7893):

okay, this is how u'd set up the matrix (im 90% sure): [-4 1] [1 -3] and now just get it's inverse.. for that, just multiply it by the identity matrix, so your inverse would be: [-4 1] [1 0] [1 -3] [0 1]

OpenStudy (accessdenied):

multiplying by the identity gets you the same matrix tho... I think the inverse is defined by [a b] -1 [d -b] [c d] = [-c a] * determinant of the matrix

OpenStudy (bahrom7893):

wait.. im sure u had to multiply it by some other matrix..

OpenStudy (bahrom7893):

but u're right, i was being an idiot

OpenStudy (accessdenied):

when you multiply the inverse by the right side of that equation i posted, that gets you the answer... but I didn't think that was the inverse, and that's what caught me up here. :P

OpenStudy (bahrom7893):

lol good answer mate

OpenStudy (accessdenied):

[-4 1] -1 [ 1 -3] [-3 -1] [-1 -4] * det(original matrix) det(original matrix) = -4*-3 - 1 = 12 - 1 = 11 [-3 -1] [-1 -4] * 11 [-33 -11] = [-11 -44] is what i get for that inverse